OpenStudy (anonymous):

find the point on the parabola 2y=x^2 that is closest to the point (-4,0)

OpenStudy (anonymous):

a point on the curve looks like \((x, \frac{x^2}{2})\)

OpenStudy (anonymous):

and the square of the distance between \((x, \frac{x^2}{2})\) and \((-4,0)\) is \[d^2=(x+4)^2+(\frac{x^2}{2})^2\] minimize that by taking the derivative, finding the critical points, etc

OpenStudy (anonymous):

thank youu!

OpenStudy (anonymous):

yw